OpenStudy (anonymous):

Solve for x. Show all your work and check you answers. 1 over x^2-7x+10=x over x-5+1over x-2 The *overs* would be a fraction but I haven't quite figured out how to make them fractions yet so sorry about that

OpenStudy (anonymous):

\[\frac{ 1 }{ x ^{2}-7x+10 }=\frac{ x }{ x-5 }+\frac{ 1 }{ x-2 }\]

OpenStudy (anonymous):

1/(x^2-7x+10)=[x/*(x-5)]+[1/(x-2)]

OpenStudy (anonymous):

\[\frac{ 1 }{ (x-5)(x-2) } = \frac{ x }{ x-5 } + \frac{ 1 }{ x-2 }\]

OpenStudy (anonymous):

to just the right side of the equation \[[\frac{ x }{ x-5 }\times\frac{ x-2 }{ x-2 }]+[\frac{ 1 }{ x-2 }\times\frac{ x-5 }{ x-5 }]+\]

OpenStudy (anonymous):

\[= \frac{ x(x-2)+1(x-5) }{ (x-2)(x-5) }\]

OpenStudy (anonymous):

\[=\frac{ (x^2-2x+x-5) }{ (x-2)(x-5 }\]

OpenStudy (anonymous):

\[=\frac{ (x^2-x-5) }{ (x-2)(x-5) }\]

OpenStudy (anonymous):

\[\frac{ 1 }{ (x-5)(x-2) }=\frac{ x^2-x-5 }{ (x-2)(x-5) }\]

OpenStudy (anonymous):

cross multiply top left with bottom right = top right with bottom left

OpenStudy (anonymous):

\[1\times(x-2)(x-5)=(x^2-x-5)(x-5)(x-2)\]

OpenStudy (anonymous):

add them together set them on one side where other side is =0 and use b square all over two a stuff and you get x=-2 and x=3

OpenStudy (anonymous):

-b +/- square root of b squared-4(a)(c)divided by 2a

OpenStudy (anonymous):

example:1x^2+3x+4=0 a=1 b=3 c=4

OpenStudy (jhannybean):

\[\small\begin{align} \frac{ 1 }{ x ^{2}-7x+10 }=\frac{ x }{ x-5 }+\frac{ 1 }{ x-2 } \ &: ~ \frac{ 1 }{ (x-5)(x-2) } = \frac{ x }{ x-5 } + \frac{ 1 }{ x-2 } \\&: ~ \frac{1}{(x-5)(x-2)}=\frac{x(x-2)}{(x-5)(x-2)}+\frac{1(x-5)}{(x-5)(x-2)} \\&: ~ (x-5)(x-2)\left[\frac{1}{(x-5)(x-2)}=\frac{x(x-2)}{(x-5)(x-2)}+ \\ \frac{1(x-5)}{(x-5)(x-2)}\right]\\ \\&: ~ 1=x(x-2)+1(x-5) \\&: ~ 1 = x^2-2x +x-5 \\&: ~ 1=x^2 -x-5 \\&:~ 0=x^2 -x-6 \end{align}\]Find two numbers that multiply to give you -6 but add to give you -1, write it out in factored form, and you're done.